# Behind Closed Words: Creating and Investigating the forePLay Annotated Dataset for Polish Erotic Discourse Warning: This paper contains examples of sexually explicit material

### v2

@@ Abstract
The surge in online content has created an urgent demand for robust detection systems, especially in non-English contexts where current tools demonstrate significant limitations. We present forePLay, a novel Polish language dataset for erotic content detection, featuring over 24k annotated sentences with a multidimensional taxonomy encompassing ambiguity, violence, and social unacceptability dimensions. Our comprehensive evaluation demonstrates that specialized Polish language models achieve superior performance compared to multilingual alternatives, with transformer-based architectures showing particular strength in handling imbalanced categories. The dataset and accompanying analysis establish essential frameworks for developing linguistically-aware content moderation systems, while highlighting critical considerations for extending such capabilities to morphologically complex languages. [^1] The forePLay dataset and associated resources will be made publicly available for research purposes upon publication, in accordance with data sharing regulations.
@@ 1 Introduction
Securing large language models (LLMs) against the generation of harmful content is another critical concern Mahomed et al. (2024); Wu et al. (2024). While traditional content moderation tools are useful for user-generated content, they are less effective as input/output guardrails for LLMs due to their inability to adapt to new policies and distinguish between user-generated and AI-generated content Inan et al. (2023). As a result, there is a growing need for datasets and specialised tools that can be integrated with LLMs to prevent the generation of explicit or harmful material, while maintaining their functionality across different applications.
While much of the previous work has focused on detecting toxicity, abusiveness, offensive language, or hate speech, Pavlopoulos et al. (2020); Gehman et al. (2020); MacAvaney et al. (2019) these categories do not encompass the full spectrum of undesired content Markov et al. (2023). Erotic material, in particular, poses a significant risk, especially to minors. Detecting such content is essential not only for creating a safer online environment but also for adhering to ethical standards and legal regulations.
@@
To address these gaps, we introduce forePLay, the first Polish language manually annotated dataset of erotic content. In contrast to prior efforts, we present a novel multidimensional taxonomy applied to a large Polish language corpus ($n=24,768$), incorporating fine-grained annotations across ambiguity, violence, and social unacceptability dimensions, while ensuring representation of LGBT-specific content from both literary and web-based sources. The article provides a detailed description of the annotation process and discusses the challenges encountered during annotation, in particular human label variation. A major contribution is also the comprehensive evaluation of erotic content detection models, examining specialised Polish transformer-based models (HerBERT Mroczkowski et al. (2021) and Polish RoBERTa Dadas (2023)) and Polish-specific LLMs such as the PLLuM family and Bielik Ociepa et al. (2024). We also compare these models with state-of-the-art multilingual and general-purpose models, including both open-source (Mixtral Jiang et al. (2024), Llama 3.1 Dubey et al. (2024), and C4AI Command-R CohereForAI (2024)) and commercial solutions (GPT-4o Hurst et al. (2024)).
@@ 3.1 Data Collection
We present a large-scale Polish language corpus ($n=24,768$ sentences) constructed through systematic sampling from two distinct sources: (i) user-generated content from online fiction stories repositories, and (ii) Polish literary works, including Polish translations of world literature. The corpus includes demographically diverse content from 905 unique text units, with a significant manually curated LGBTQ+ representation. Document-level context is preserved through unique file identifiers. While direct source attribution via URLs is omitted due to the volatility of online content, our sampling methodology ensures comprehensive coverage of various linguistic registers and social contexts.
Given the potential application of models in supporting moderation, we focused on internet language and amateur writing, which is predominantly anonymous. Consequently, a significant portion of the sentences (69%) originates from online stories. While the majority of these stories are erotic and non-professional, we included non-erotic stories in a 1:4 ratio to enhance diversity and reduce genre-specific biases. All non-professional stories were scraped from publicly available websites, see Appendix A for details.
@@ 4 Annotation Process
To maintain inclusiveness and diversity in the annotation process, we assembled a gender-balanced team of annotators, comprising 3 male and 3 female individuals, aged 20-40. To minimize potential bias, the annotators were not provided with metadata that could link specific samples to genre types. Additionally, the samples from each genre type were distributed evenly among all annotators. We included annotators with backgrounds in linguistics, literature, or related fields to ensure sensitivity to language nuances. Each sentence was independently assessed by three annotators, with final labels determined by majority vote. In cases where all three labels differed, a superannotator made the final decision. This occurred in 830 out of $24,768$ samples, representing 3.35% of the total number of observations.
@@ 4.1 Annotation Scheme
Based on the expert knowledge of the content, a set of five possible exclusive labels to annotate the samples was designed. This included the following categories: erotic (e), ambigous (a), violence-related (v), socially unacceptable behaviors (u), and neutral (n).
@@
The annotation guidelines, developed by literary scholars and linguists based on domain expertise and material analysis, are described in detail in B.
@@
Table 2: Distribution of Label Categories Across Dataset Subcorpora
@@
Table 3: Examples of annotations in the dataset
Original sentence in Polish English translation Label
Był cały twardy. He was all hard. ambiguous
Ten gad ją zdradzał! This rat was cheating on her! neutral
@@ 4.2 Annotation Quality
Human label variation (HLV) Plank (2022), inherently present in our task of annotating erotic sentences, raised several questions regarding the dataset design (see the 7 Section). To preserve the authenticity of the annotation process, we chose not to intervene beyond the initial training phase. For the dataset’s initial release, we opted to follow the conventional practice of data aggregation featuring a ground truth.
@@
Agreement in most pairs was consistent, ranging from 0.66 to 0.71, regardless of annotators’ gender. However, pairs involving annotator F1 presented a notable deviation, as F1 emerged as an outlier, largely due to an excessive use of the ambiguous label.
In addition to pairwise agreement, the overall consistency of the entire sample was assessed using Krippendorff’s Alpha. The value was quite low (0.387). However, after excluding the F1 annotation, which significantly deviated from the other annotators, the value increased to 0.716, indicating a satisfactory level of agreement in the annotation process.
Table 4: Cohen’s Kappa values for each annotator pair across all categories and for each individual category, broken down by gender (F - female, M - male).
@@ 6 Results
Table 6 highlights the F1 performance metrics of the HerBERT and RoBERTa models across classification tasks involving different dataset configurations. In binary classification scenarios, both models exhibit robust performance. However, as the number of classes increases, a decline in F1 scores is observed. Notably, RoBERTa, particularly in its Base configuration, demonstrates superior performance relative to HerBERT, with pronounced advantages evident in tasks on the Basic and Extended datasets.
Table 5: F1 scores for various models across different dataset configurations and shot numbers.
Model Base Core Extended Full
@@
PLLuM-Mistral-12B (SFT) 0.946 0.792 0.458 0.488
In the case of LLMs, models from the PLLuM family consistently demonstrate strong performance across varying label counts and shot configurations, with particular emphasis on the PLLuM-Mixtral-8x7B, which achieves outstanding F1 scores across a wide range of label counts, notably reaching 0.921 for Basic data and 0.670 for Core data in the 5-shot setting. Additionally Llama-3.1-8B-PLLuM (SFT) and PLLuM-Mistral-12B (SFT) show significant improvements when subjected to supervised fine-tuning, resulting in enhanced performance, with the highest F1 scores observed across various label counts.
Table 6: F1 scores for HerBERT and RoBERTa models across different dataset versions.
Dataset HerBERT RoBERTa
Base Large Base Large
Base 0.929 0.939 0.944 0.943
Core 0.702 0.738 0.738 0.748
Extended 0.693 0.746 0.704 0.734
Full 0.632 0.648 0.707 0.664
Fine-tuned LLMs exhibit performance levels comparable to traditional transformer-based architectures when evaluated on datasets labeled as Base and Core, where all categories are sufficiently represented. This suggests that in scenarios characterized by balanced class distributions, where each class is supported by an adequate number of training examples, the performance differential between these methodological approaches is minimal. Conversely, for datasets labeled as Extended and Core, which are marked by pronounced class imbalance, models built upon the BERT architecture begin to demonstrate a distinct advantage. These models excel in effectively addressing the challenges posed by imbalanced data distributions, exhibiting superior capability in identifying underrepresented categories. This highlights their enhanced adaptability and robustness in handling more complex scenarios involving skewed training data distributions.
@@
Table 7: Number of unidentified responses or refusals.
Model Base Core Extended Full Total
@@ 7 Discussion
Our experimental results demonstrate clear performance patterns across model architectures and linguistic contexts, with specialized Polish language models consistently outperforming multilingual alternatives in erotic content detection. The encoder-based architectures, particularly RoBERTa, achieve robust F1 scores in binary classification ($0.929-0.944$), though performance degrades with increasing categorical complexity (dropping to $\sim 0.66$ for five-class classification). This degradation pattern reveals fundamental challenges in maintaining discriminative power across finer-grained categorical distinctions, particularly in cases requiring subtle cultural and contextual understanding.
The PLLuM family of models exhibits exceptional few-shot learning capabilities, with PLLuM-Mixtral-8x7B achieving F1 scores of $0.921$ in 5-shot settings for binary classification, while supervised fine-tuning of PLLuM-Mistral-12B yields the highest overall performance (F1=$0.946$). However, the diminishing returns observed with increasing model size, coupled with the persistent challenge of handling ambiguous content in imbalanced datasets, suggests that architectural sophistication alone may not overcome the fundamental challenges of erotic content detection in morphologically complex languages. These findings underscore the importance of language-specific model development for sensitive content detection tasks, rather than relying solely on increasingly large multilingual models.
@@ 8 Conclusions
9 Limitations
@@ Appendix B Annotation Guidelines Summary
Annotators are instructed to carefully read each sentence and assign one final label to it. They are also encouraged to pause or discontinue the task at any point if they feel uncomfortable with the content being evaluated.
Each sample should be assigned with one of the following labels:
Label e (for erotic): choose this label if a given sentence describes sexual activities and desires, advanced flirting with evident erotic undertones, references to past sexual activities, or explicit sexual fantasies. Please note that mere mentions of genital terms or publicly accepted romantic behaviors or neutral discussions of erotic or sexual topics (e.g., from the psychological perspective) are excluded from this category.
Label v (for violence-related): choose this label if a given sample depicts sexual behaviors widely considered illegal, taboo, or violative of societal norms. Examples can include references to zoophilia, necrophilia, pedophilia, incest, or other sexual deviations. Please note that a sentence can include violence which is not sex-related. Such instances are excluded from this category.
Label u (for socially unacceptable behaviors): choose this label if a given sentence describes sexual behaviors considered illegal, generally taboo, or violative of social norms, such as zoophilia, necrophilia, pedophilia, incest, and other sexual deviations. Please note that if you classify a sample as both violence related and featuring socially unacceptable behaviors, the priority should be given to the latter. Such instances should be label as socially unacceptable behaviors.
Label a (for ambiguous): choose this label to sentences that may evoke some degree of erotic connotation but whose interpretation depends heavily on the context. If the sentence can reasonably be interpreted as non-erotic in a neutral context, it should be labeled as ambiguous.
Label n (for neutral): choose this label if a given sentence does not meet the criteria of the above mentioned sex-related categories. This typically includes non-sexual content on various topics, as well as general statements about human sexuality or intimate anatomy that do not carry erotic connotations.

### v3

@@ Abstract
The surge in online content has created an urgent demand for robust detection systems, especially in non-English contexts where current tools demonstrate significant limitations. We introduce forePLay, a novel Polish-language dataset for erotic content detection, comprising over 24,000 annotated sentences. The dataset features a multidimensional taxonomy that captures ambiguity, violence, and socially unacceptable behaviors. Our comprehensive evaluation demonstrates that specialized Polish language models achieve superior performance compared to multilingual alternatives, with transformer-based architectures showing particular strength in handling imbalanced categories. The dataset and accompanying analysis establish essential frameworks for developing linguistically-aware content moderation systems, while highlighting critical considerations for extending such capabilities to morphologically complex languages. [^1] Dataset link: https://github.com/ZILiAT-NASK/ForePLay
@@ 1 Introduction
Securing large language models (LLMs) against the generation of harmful content is another critical concern Mahomed et al. (2024); Wu et al. (2024). While traditional content moderation tools are useful for user-generated content, they are less effective as input/output guardrails for LLMs due to their inability to adapt to new policies and distinguish between user-generated and AI-generated content Inan et al. (2023). As a result, there is a growing need for datasets and specialized tools that can be integrated with LLMs to prevent the generation of explicit or harmful material, while maintaining their functionality across different applications.
While much of the previous work has focused on detecting toxicity, abusiveness, offensive language, or hate speech Pavlopoulos et al. (2020); Gehman et al. (2020); MacAvaney et al. (2019), these categories do not encompass the full spectrum of undesired content Markov et al. (2023). Erotic material, in particular, poses a significant risk, especially to minors. Detecting such content is essential not only for creating a safer online environment but also for upholding ethical standards. Although the Digital Services Act (DSA) primarily addresses non-consensual pornography and age-verification requirements, many online platforms, particularly those targeting minors, enforce stricter moderation policies that extend beyond legal obligations. Some platforms may prohibit explicit erotica or any 18+ material, even when it does not violate the law. One application of this research is to support such platforms in enforcing their internal guidelines, which vary depending on the service and its target audience.
@@
To address these gaps, we introduce forePLay, the first Polish language manually annotated dataset of erotic content. In contrast to prior efforts, we present a novel multidimensional taxonomy applied to a large Polish language corpus ($n=24,768$), incorporating fine-grained annotations across ambiguity, violence, and social unacceptability dimensions, while ensuring representation of LGBT-specific content from both literary and web-based sources. The corpus is constructed from user-generated content sourced from online fiction repositories and professionally published Polish literary works, offering diverse linguistic and stylistic coverage. The article provides a detailed description of the annotation process and discusses the challenges encountered during annotation, in particular human label variation. A major contribution is also the comprehensive evaluation of erotic content detection models, examining specialized Polish transformer-based models (HerBERT Mroczkowski et al. (2021) and Polish RoBERTa Dadas (2023)) and Polish-specific LLMs such as the PLLuM family [^2] For more information, see https://pllum.org.pl/ and https://huggingface.co/CYFRAGOVPL and Bielik Ociepa et al. (2024). We also compare these models with state-of-the-art multilingual and general-purpose models, including both open-source (Mixtral Jiang et al. (2024), Llama 3.1 Dubey et al. (2024), and C4AI Command-R CohereForAI (2024)) and commercial solutions (GPT-4o Hurst et al. (2024)).
@@ 3.1 Data Collection
We constructed a large-scale Polish language corpus ($n=24{,}768$ sentences) through systematic sampling from two distinct sources: (i) user-generated content from online fiction repositories and (ii) Polish literary works, including translations of world literature. The data collection was carried out by an interdisciplinary team consisting of literary scholars with extensive expertise in contemporary Polish literature, sociologists, NLP specialists, and IT engineers. The corpus comprises demographically diverse material drawn from 905 unique text units and includes a substantial, manually curated representation of LGBTQ+ narratives. This representation was ensured by selecting queer-themed literary texts—primarily original Polish works, supplemented by a smaller number of translated pieces—based on scholarly research and expertise in Polish queer literature. Document-level context is preserved through unique file identifiers. While direct URLs are omitted due to the volatility of online sources, our sampling methodology ensures broad coverage across linguistic registers and social contexts.
Given the potential application of models in supporting moderation, we focused on internet language and amateur writing, which is predominantly anonymous. Consequently, a significant portion of the sentences (69%) originates from online stories. While the majority of these stories are erotic and non-professional, we included non-erotic stories in a 1:4 ratio to enhance diversity and reduce genre-specific biases. To filter erotic content, we relied on tags and category labels related to erotica provided by the online repositories. All non-professional stories were scraped from publicly available websites, see Appendix A for details.
@@ 4 Annotation Process
To assemble the annotation team, three female annotators were allocated from our internal organization. To ensure gender balance, three part-time male annotators were recruited as external contractors. All selected annotators underwent training and guidance to maintain annotation consistency and quality throughout the project.
This process resulted in a gender-balanced team of six annotators (3 male, 3 female), aged 20–40. We included annotators with backgrounds in linguistics, literature, or related fields to ensure sensitivity to language nuances. To minimize potential bias, the annotators were not provided with metadata that could link specific samples to genre types. Furthermore, samples from each genre type were distributed evenly among all annotators. Each sentence was independently assessed by three annotators. Final labels were determined by majority vote. In cases where all three labels differed ($830$ out of $24,768$ samples, representing $3.35\%$ of the total observations), a superannotator made the final decision. The superannotator, an NLP specialist and senior member of the annotation team, was responsible for overseeing consistency and quality throughout the annotation process. This role encompassed resolving disagreements among annotators, providing additional guidance, and conducting spot checks to ensure adherence to the defined guidelines.
@@ 4.1 Annotation Scheme
Based on the expert knowledge of the content, a set of five possible exclusive labels to annotate the samples was designed. This included the following categories: erotic (e), ambiguous (a), violence-related (v), socially unacceptable behaviors (u), and neutral (n). Some examples of the annotated text and its English translation are presented in Table 2.
Original sentence in Polish English translation Label
Był cały twardy. He was all hard. ambiguous
Ten gad ją zdradzał! This rat was cheating on her! neutral
Table 2: Examples of annotations in the dataset
@@
The annotation guidelines, developed by literary scholars and linguists based on domain expertise and material analysis, are described in detail in Appendix B.
@@
Table 3: Distribution of Label Categories Across Dataset Subcorpora
@@ 4.2 Annotation Quality
Human label variation (HLV) Plank (2022), inherently present in our task of annotating erotic sentences, raised several questions regarding the dataset design (see the Section 7). To preserve the authenticity of the annotation process, we chose not to intervene beyond the initial training phase. For the dataset’s initial release, we opted to follow the conventional practice of data aggregation featuring a ground truth.
@@
Agreement in most pairs was consistent, ranging from 0.66 to 0.71, regardless of annotators’ gender. However, pairs involving annotator Fem1 presented a notable deviation, as Fem1 emerged as an outlier, largely due to an excessive use of the ambiguous label.
In addition to pairwise agreement, the overall consistency of the entire sample was assessed using Krippendorff’s Alpha. The value was quite low (0.387). However, after excluding the Fem1 annotation, which significantly deviated from the other annotators, the value increased to 0.716, indicating a satisfactory level of agreement in the annotation process.
@@
Table 4: Cohen’s Kappa values for each annotator pair across all categories and for each individual category, broken down by gender (Fem – female, M – male).
@@ 6 Results
Table 5 highlights the macro-F1 performance metrics of the HerBERT and RoBERTa models across classification tasks involving different dataset configurations. In binary classification scenarios, both models exhibit robust performance. However, as the number of classes increases, a decline in macro-F1 scores is observed. Notably, RoBERTa, particularly in its Base configuration, demonstrates superior performance relative to HerBERT, with pronounced advantages evident in tasks on the Basic and Extended datasets.
Dataset HerBERT RoBERTa
Base Large Base Large
Basic 0.929 0.939 0.944 0.943
Core 0.702 0.738 0.738 0.748
Extended 0.693 0.746 0.704 0.734
Full 0.632 0.648 0.707 0.664
Table 5: Macro-F1 scores for HerBERT and RoBERTa models across different dataset versions.
Model Basic Core Extended Full
@@
Table 6: Macro-F1 scores for various models across different dataset configurations and shot numbers.
In the case of LLMs, models from the PLLuM family consistently demonstrate strong performance across varying label counts and shot configurations, with particular emphasis on the PLLuM-Mixtral-8x7B, which achieves outstanding macro-F1 scores across a wide range of label counts, notably reaching 0.921 for Basic data and 0.670 for Core data in the 5-shot setting. Additionally Llama-3.1-8B-PLLuM (SFT) and PLLuM-Mistral-12B (SFT) show significant improvements when subjected to supervised fine-tuning, resulting in enhanced performance, with the highest macro-F1 scores observed across various label counts.
Fine-tuned LLMs exhibit performance levels comparable to traditional transformer-based architectures when evaluated on datasets labeled as Basic and Core, where all categories are sufficiently represented. This suggests that in scenarios characterized by balanced class distributions, where each class is supported by an adequate number of training examples, the performance differential between these methodological approaches is minimal. Conversely, for datasets labeled as Extended and Core, which are marked by pronounced class imbalance, models built upon the BERT architecture begin to demonstrate a distinct advantage. These models excel in effectively addressing the challenges posed by imbalanced data distributions, exhibiting superior capability in identifying underrepresented categories. This highlights their enhanced adaptability and robustness in handling more complex scenarios involving skewed training data distributions.
@@
Model Basic Core Extended Full Total
@@ 7 Discussion
Our experimental results demonstrate clear performance patterns across model architectures and linguistic contexts, with specialized Polish language models consistently outperforming multilingual alternatives in erotic content detection. The encoder-based architectures, particularly RoBERTa, achieve robust macro-F1 scores in binary classification ($0.929-0.944$), though performance degrades with increasing categorical complexity (dropping to $\sim 0.66$ for five-class classification). This degradation pattern reveals fundamental challenges in maintaining discriminative power across finer-grained categorical distinctions, particularly in cases requiring subtle cultural and contextual understanding.
The PLLuM family of models exhibits exceptional few-shot learning capabilities, with PLLuM-Mixtral-8x7B achieving macro-F1 scores of $0.921$ in 5-shot settings for binary classification, while supervised fine-tuning of PLLuM-Mistral-12B yields the highest overall performance (macro-F1 = $0.946$). However, the diminishing returns observed with increasing model size, coupled with the persistent challenge of handling ambiguous content in imbalanced datasets, suggests that architectural sophistication alone may not overcome the fundamental challenges of erotic content detection in morphologically complex languages. These findings underscore the importance of language-specific model development for sensitive content detection tasks, rather than relying solely on increasingly large multilingual models.
Figure 1: Type I errors percentage across datasets and inference settings. Error bars represent 95% confidence interval.
Beyond overall performance metrics, understanding error patterns is crucial for content moderation deployment, where both false positives (over-censorship) and false negatives (missed inappropriate content) carry significant consequences. We calculated Type I error percentages as the number of false positives divided by the sum of false positives and false negatives to assess model precision across different classification scenarios. Full results regarding Type I and Type II errors for each model are presented in the Appendix D. The analysis reveals that Type I error rates decrease substantially with classification complexity—from approximately 60-80% in binary classification (Basic) to 40-50% in the full five-class scenario (Figure 1). This pattern indicates higher Type I error rates in simpler tasks and higher Type II rates in complex multi-class scenarios, highlighting the fundamental trade-off between precision and sensitivity in erotic content detection tasks.
@@ 8 Conclusions
9 Data Availability
A subset of erotic and ambiguous sentences, totaling $3{,}704$ samples, has been released on our public Github account. Due to ethical considerations, we opted to exclude the minor harmful classes of violence-related content and socially unacceptable behaviors. ForePLay Dataset Release 1.0 consists of $2{,}728$ sentences labeled as erotic, and $976$ sentences labeled as ambiguous, which corresponds to 43% and 73% of the original dataset, respectively. Neutral sentences were not included in this release, as such data is readily obtainable from publicly available corpora and may be easily used for comparative purposes in downstream tasks. This subset has undergone additional copyright verification and was made available following legal consultations. The released data and its license comply with new legal regulations that came into effect after the data collection and annotation process had been completed.
Limitations
@@ 9 Data Availability
Furthermore, a specific limitation pertaining to the evaluation phase is that the definitions of the labels were not explicitly included in the prompts provided to the models. We confirm this omission may have contributed to the lower performance of the models in comparison to the baseline.
@@
Acknowledgements
The development of the PLLuM models used in the experiments was funded by the Polish Minister of Digital Affairs under a special purpose subsidy No. 1/WI/DBiI/2023: Responsible Development of the Open Large Language Model, PLLuM (Polish Large Language Model), aimed at supporting breakthrough technologies in the public and economic sectors, including an open, Polish-language intelligent assistant for public administration clients. The work on the dataset, however, was conducted independently and was not financed by this project.
@@ Appendix B Annotation Guidelines Summary
Annotators are instructed to carefully read each sentence and assign one final label to it. Each sentence was presented in isolation, without additional context. They are also encouraged to pause or discontinue the task at any point if they feel uncomfortable with the content being evaluated.
Each sample should be exclusively assigned one of the following labels:
  • Label e (for erotic): choose this label if the sentence describes sexual activities and desires, advanced flirting with evident erotic undertones, references to past sexual activities, or explicit sexual fantasies.
    Please note that mere mentions of genital terms or publicly accepted romantic behaviors or neutral discussions of erotic or sexual topics (e.g., from the psychological perspective) are excluded from this category.
  • Label v (for violence-related): choose this label if the sentence depicts explicit sexual harassment, rape, lack of consent, or any other sort of non-consensual violence which concerns sexual intentions or activities.
    Please note that a sentence can include violence which is not sex-related. Such instances are excluded from this category. Consensual BDSM behaviors are likely to fall under the label erotic.
  • Label u (for socially unacceptable behaviors): choose this label if the sentence describes sexual behaviors considered illegal, generally taboo, or violative of social norms, such as zoophilia, necrophilia, pedophilia, incest, and other sexual deviations.
    Please note that if you classify a sample as both violence related and featuring socially unacceptable behaviors, the priority should be given to the latter. Such instances should be labeled as socially unacceptable behaviors.
  • Label a (for ambiguous): choose this label for sentences that may evoke some degree of erotic connotation but whose interpretation depends heavily on the context. If the sentence can reasonably be interpreted as non-erotic in a neutral context, it should be labeled as ambiguous.
  • Label n (for neutral): choose this label if the sentence does not meet the criteria of the above mentioned sex-related categories. This typically includes non-sexual content on various topics, as well as general statements about human sexuality or intimate anatomy that do not carry erotic connotations.